# MOOSE-Chem3: Toward Experiment-Guided Hypothesis Ranking via Simulated Experimental Feedback

### v2

@@ Abstract
Hypothesis ranking is a crucial component of automated scientific discovery, particularly in natural sciences where wet-lab experiments are costly and throughput-limited. Existing approaches focus on pre-experiment ranking, relying solely on large language model’s internal reasoning without incorporating empirical outcomes from experiments. We introduce the task of experiment-guided ranking, which aims to prioritize candidate hypotheses based on the results of previously tested ones. However, developing such strategies is challenging due to the impracticality of repeatedly conducting real experiments in natural science domains. To address this, we propose a simulator grounded in three domain-informed assumptions, modeling hypothesis performance as a function of similarity to a known ground truth hypothesis, perturbed by noise. We curate a dataset of 124 chemistry hypotheses with experimentally reported outcomes to validate the simulator. Building on this simulator, we develop a pseudo experiment-guided ranking method that clusters hypotheses by shared functional characteristics and prioritizes candidates based on insights derived from simulated experimental feedback. Experiments show that our method outperforms pre-experiment baselines and strong ablations. [^1] All code and data can be found in https://github.com/wanhaoliu/MOOSE-Chem3.gitBoth authors contributed equally to this work. Corresponding author.
@@ 1 Introduction
Scientific discovery plays a major role in advancing human society (Coccia, 2019). Recently, there have been promising advances in automating certain stages of the scientific process using large language models (LLMs) (Luo et al., 2025; Cambria et al., 2023).
@@
One of the most critical stages is hypothesis ranking: given a large set of automatically generated hypotheses, which one should be tested in a real experiment first? This question is particularly important in natural science domains, where experiments are costly and resource-constrained, necessitating efficient prioritization strategies to minimize experimental effort.
Previous methods for hypothesis ranking (Yang et al., 2024b; Si et al., 2024) primarily rely on evaluations based solely on LLMs’ internal reasoning, without incorporating any empirical feedback from experiments. We refer to this approach as pre-experiment ranking, as hypotheses are prioritized before any experimental evidence is gathered.
In contrast, we propose a new task: experiment-guided ranking, which focuses on dynamically prioritizing hypotheses by leveraging feedback from sequentially performed experiments. Rather than conducting all experiments upfront, this approach iteratively updates the ranking based on available experimental results, aiming to accelerate the discovery of promising hypotheses while minimizing the total number of experiments required. However, developing strategies for experiment-guided ranking in natural science domains such as chemistry is challenging, as it is impractical to rely on real laboratories to repeatedly conduct experiments. In other words, the lack of scalable access to meaningful experimental feedback remains a key barrier to researching experiment-guided ranking strategies.
Despite the challenges of obtaining real experimental feedback, we posit that simulating such feedback is feasible under three foundational assumptions. To illustrate these, consider a latent space where the $x$-axis (potentially multidimensional) parameterizes candidate hypotheses, such that each coordinate corresponds to a distinct hypothesis variant, and the $y$-axis denotes the associated experimental feedback (e.g., performance). Assumption 1 (A1) posits that within any sufficiently local neighborhood of the hypothesis space, there exists at most one dominant optimum, corresponding to a ground truth hypothesis (e.g., reported in the literature). Assumption 2 (A2) states that hypotheses closer to this dominant maximum are more likely to yield more competitive experimental feedback. Assumption 3 (A3) states that real experimental feedback can be viewed as idealized feedback—defined by A1 and A2—perturbed by an unknown deviation term due to the imperfect representation of hypothesis closeness in the hypothesis space.
Specifically, the ground truth hypothesis is treated as the local optimum in the hypothesis space, and the performance of neighboring hypotheses is modeled as a function of their similarity to this optimum. Since real-world representations of hypothesis similarity are inherently imperfect, this relationship is subject to systematic deviations, whose effects on the simulator’s fidelity are analyzed in this work.
The experiment-guided ranking task with real and simulated experiment feedback can be described by Figure 1. The primary goal of the simulator is to enable systematic research on experiment-guided ranking strategies by providing accessible and high-fidelity approximations of experimental feedback, which are otherwise prohibitively costly or unavailable. Ultimately, the aim is to deploy these strategies in real experimental settings to reduce the overall experimental costs.
(a) Real experiment feedback.
(b) Simulated experiment feedback.
Figure 1: Experiment-guided hypothesis ranking using real and simulated feedback. $A1$, $A2$, and $A3$ illustrate our three foundational assumptions in a concise manner (introduced in § 2.1.1).
Guided by these assumptions and insights from chemistry experts, we construct a simulator that closely approximates real wet-lab experimental outcomes. To evaluate its fidelity, we curate a dataset of 30 groups of chemistry hypotheses, each consisting of 3–6 related hypotheses along with their experimentally reported performance, sourced from published literature (in total 124 <hypothesis, performance> pairs). Our simulator outperforms strong baselines, including widely used similarity-based evaluation metrics for hypothesis comparison (Yang et al., 2024b). Building on this foundation, we propose a new task: developing more accurate simulators for experimental feedback. Using the simulator to approximate experimental outcomes, we develop an pseudo experiment-guided ranking method leveraging functional clustering of hypotheses. Clustering enables effective transfer of insights from previously tested hypotheses to untested ones sharing similar functional elements, rather than evaluating each hypothesis in isolation.
Specifically, hypotheses containing elements with similar functional relevance—regardless of whether identical or distinct—are grouped together, allowing hypotheses to belong to multiple clusters. Our method prioritizes clusters based on accumulated experimental feedback and subsequently selects the most promising hypothesis within each. Experiments demonstrate that our approach outperforms both pre-experiment ranking methods and strong ablation variants. Overall, the contributions of this paper are:
@@
  • We formalize the task of experiment-guided ranking and highlight a key challenge in the natural sciences: the lack of scalable access to wet-lab experimental feedback. To address this, we propose the use of simulators and release a curated dataset of 124 chemical hypotheses with annotated performance collected from the literature.
@@
  • We introduce three foundational assumptions for simulating experimental feedback, provide a mathematical formalization of the simulation process, and construct a high-fidelity simulator that approximates real wet-lab outcomes under these assumptions.
  • We develop a clustering-based pseudo experiment-guided ranking method that leverages simulated feedback and structural similarities among hypotheses. Experimental results show that our method outperforms both pre-experiment baselines and strong ablation variants.
@@ 2 Methodology of Simulator Construction
2.1 Foundational Assumptions and Formalization
Our simulator construction is guided by three foundational assumptions derived from expert consultations in the chemistry domain. These assumptions provide a principled basis for modeling simulated experimental outcomes of untested chemical hypotheses, enabling systematic investigation of experiment-guided ranking strategies.
2.1.1 Foundational Assumptions
(a) Idealized performance landscape ($A1$ + $A2$)).
(b) Realistic performance landscape ($A1$ + $A2$ + $A3$).
(c) Deviations from imperfect closeness estimation ($A3$).
Figure 2: Illustration of the three fundamental assumptions for simulator construction.
We posit that real experimental feedback within a hypothesis space can be simulated under the following assumptions:
@@ 2.1.1 Foundational Assumptions
    (A1) Within any sufficiently local neighborhood of the hypothesis space, there exists at most one dominant optimum, corresponding to a ground truth hypothesis.
@@
    (A2) Among hypotheses in the vicinity of a dominant optimum, those that are closer to it are more likely to yield better experimental feedback.
@@
    (A3) Real experimental outcomes deviate from the idealized structure described in A1 and A2 due to the imperfect representation of hypothesis closeness in the hypothesis space.
Figure 2 visually illustrates these assumptions. In the ideal scenario (Figure 2(a)), hypotheses are embedded in a latent hypothesis space such that the Euclidean distance (“closeness”) between a hypothesis and the dominant optimum hypothesis accurately reflects similarity in terms of how they perform on a research question, creating a smooth, unimodal performance landscape.
However, practical scenarios differ substantially, since the distance (“closeness”) between hypotheses—whether assessed by scientists or LLMs—may not accurately reflect functional similarity. For example, a chemical hypothesis may include a useful functional component whose contribution is not fully recognized, causing it to be placed farther from the dominant peak than it should be—resulting in a spurious secondary peak. Conversely, a suboptimal hypothesis may appear closer to the dominant peak than warranted, forming a local valley. These distortions result in a performance landscape such as that in Figure 2(b), with unexpected secondary peaks and valleys. Figure 2(c) further isolates this deviation, representing the discrepancy between oracle and practical understandings of closeness.
We now formalize these assumptions by defining a mathematical model that makes the relationship between hypothesis embeddings, similarity, and performance explicit.
@@ 2.1.2 Mathematical Formulation
$f(h,h^{*};q,\phi^{*}(\cdot))=\frac{1}{(2\pi\sigma^{2})^{d/2}}\exp\left(-\frac{\|\phi^{*}(h\mid q)-\phi^{*}(h^{*}\mid q)\|^{2}}{2\sigma^{2}}\right),$ (1)
@@
We model the idealized performance surface as a Gaussian-like function centered at $\phi^{*}(h^{*}\mid q)$, yielding a strictly unimodal landscape that decays smoothly with increasing distance from the optimum $h^{*}$ (Figure 2(a)). While the true performance landscape in chemical space may not be strictly Gaussian, the isotropic Gaussian form serves as a tractable and interpretable approximation in the latent space. This modeling choice directly reflects Assumptions A1 and A2.
@@
where $\phi(\cdot\mid q)$ is a practical embedding function that maps each hypothesis $h$ into (somewhat distorted) positions in the latent hypothesis space for a research question $q$, and $\epsilon(h\mid q)$ represents a systematic correction term that accounts for the discrepancy between oracle embedding $\phi^{*}(\cdot\mid q)$ and the practical embedding $\phi(\cdot\mid q)$ under the context of $q$. As a result, the practical embedding $\tilde{\mathcal{H}}$ introduces systematic distortions in the latent space, leading to spurious local optima or unexpected valleys—effectively transforming the unimodal ideal surface into a noisier, multimodal one (Figure 2(b)).
@@
2.2 A Practical Implementation of $\phi(\cdot)$ with Chemistry Prior Knowledge
@@
As discussed in § 2.1, the core objective of the simulator is to construct an embedding function $\phi(\cdot)$ that maps each hypothesis $h$ into a latent space such that distances in this space reflect meaningful functional differences. Through extensive discussions with chemistry PhD students, we observe that a chemical hypothesis succeeds in addressing a research question primarily due to its underlying reaction mechanisms.
Specifically, an effective hypothesis typically comprises a set of chemically meaningful components—each contributing to distinct yet complementary sub-mechanisms—which together enable the overall reaction to fulfill its intended function. The specific prompts and examples for extracting key chemical components and inferring mechanisms are provided in § A.
@@
The simulation begins by decomposing both the candidate and ground truth hypotheses into a set of key chemical components, and identifying the reaction mechanism associated with each component in the context of the research question. The decomposition of $h^{*}$ is performed first, serving as a reference. These reference components and mechanisms guide the decomposition of $h$, ensuring alignment in both granularity and mechanistic interpretation.
@@
$S(h\mid q;h^{*})=\left(\prod_{i\in\mathcal{C}}\mathbf{1}_{{s_{i}>0}}\right)\cdot\left(\sum_{i=1}^{K}w_{i}\cdot s_{i}\right),\quad\text{where}\quad\sum_{i=1}^{K}w_{i}=1$ (3)
@@ 3 Methodology of Experiment-Guided Ranking
3.1 Problem Formulation
Given a research question $q$, a set of candidate hypotheses $\mathcal{H}$ is formed by selecting hypotheses generated by existing scientific discovery systems (Yang et al., 2024b) and ground-truth hypotheses from top-tier chemistry journals reporting high-quality lab experiments. The goal of experiment-guided ranking is to identify the optimal hypothesis $h^{*}\in\mathcal{H}$ with the highest experimentally measured performance using an experiment executor $E$.
Formally, we define the experiment executor as a function:

### v3

@@ Abstract
Hypothesis ranking is a crucial component of automated scientific discovery, particularly in natural sciences where wet-lab experiments are costly and throughput-limited. Existing approaches focus on pre-experiment ranking, relying solely on a language model’s internal reasoning without incorporating empirical outcomes. We introduce the task of experiment-guided ranking, which prioritizes hypotheses based on feedback from previously tested ones. However, developing such strategies in natural science domains is challenging due to the impractical requirement of repeatedly conducting real experiments. To address this, we revisit the core purpose of real experiments: to provide feedback on both the groundtruth hypothesis and the surrounding hypotheses that form the path toward it. This motivates our alternative: a simulator grounded in three domain-informed conceptual foundations, modeling hypothesis performance as a function of similarity to a known ground truth, perturbed by noise. While the groundtruth is pre-specified, it remains hidden from the ranking agent, enabling faithful evaluation of policies that navigate toward it. Validated against 124 hypotheses with experimentally reported outcomes, the simulator approximates real experimental results with consistent trend alignment. Though not perfectly accurate, its deviations resemble wet-lab noise and can foster more robust ranking strategies. We formulate experiment-guided ranking as a sequential decision-making problem and propose an in-context reinforcement learning (ICRL) framework. Within this framework, we introduce an LLM-based agentic policy that decomposes hypotheses into functional elements, clusters them by shared mechanistic roles, and prioritizes recombinations of promising elements based on feedback. Experiments show that our method significantly outperforms pre-experiment baselines and strong ablations. Our toolkit—comprising the simulator and ICRL framework—enables systematic research on experiment-guided ranking, with our policy serving as a strong proof of concept.  [^1] All code and data can be found in https://github.com/wanhaoliu/MOOSE-Chem3.gitBoth authors contributed equally to this work.
@@ 1 Introduction
Scientific discovery plays a foundational role in advancing human society (Coccia, 2019). Recent progress in large language models (LLMs) has sparked growing interest in automating parts of this scientific process (Luo et al., 2025). Among these, one of the most critical stages is hypothesis ranking: given a large set of automatically generated hypotheses (e.g., by AI), which one should be tested in a real experiment first? This question is particularly important in natural science domains, where wet-lab experiments are costly and throughput-limited, requiring prioritization strategies that maximize discovery efficiency under strict experimental budgets.
Existing work on hypothesis ranking (Yang et al., 2025; Si et al., 2024) primarily relies on evaluations based solely on a language model’s internal reasoning, without incorporating any empirical feedback. We refer to this as pre-experiment ranking. While efficient, this approach overlooks the iterative, feedback-driven nature of real-world experimentation.
In contrast, we introduce the task of experiment-guided ranking, which prioritizes hypotheses for the next round of experimentation based on outcomes from previously tested ones. Rather than evaluating all candidates upfront, this approach dynamically adjusts prioritization as new experimental results become available. However, in natural science domains such as chemistry, materials science, and biology, conducting iterative experiments at scale—as required by experiment-guided ranking—is often infeasible due to the high cost, long duration, and limited throughput of real-world experimentation. This lack of scalable feedback limits progress in developing and evaluating experiment-guided ranking strategies.
To address this challenge, we revisit the core purpose of real experiments: not only to validate a ground-truth hypothesis, but also to provide feedback on nearby hypotheses that form the path toward it. This motivates our alternative: a simulator that approximates experimental feedback in a local neighborhood of hypothesis space, enabling the development and evaluation of experiment-guided ranking strategies.
Our simulator is grounded in three conceptual foundations, reflecting the universal natural-science principle that structural similarity implies similar behavior (Callister & Rethwisch, 1999; Hansch et al., 1995; Wiley, 1986; Alberts et al., 2015). A1 (Local Optimum Assumption) states that a ground-truth hypothesis represents a dominant local optimum within its sufficiently local neighborhood. P1 (Scientific Principle) holds that greater structural or functional similarity yields more similar outcomes. D1 (Logical Deduction) follows that, because similarity representations are imperfect, the observed performance landscape deviates from the ideal implied by A1 and P1.
We formalize these conceptual foundations and construct a simulator that models hypothesis performance as a function of distance to a hidden ground-truth hypothesis. Although the ground truth is known to the simulator, it remains hidden from the ranking policy—enabling rigorous evaluation of strategies that must infer it through limited feedback. To validate the simulator, we curate a dataset of 124 hypotheses with experimentally reported outcomes from the literature. Our simulator demonstrates high trend alignment and predictive accuracy in approximating real experimental outcomes. It also outperforms strong baselines adapted from prior work (Yang et al., 2025), further supporting its utility as a research tool for developing and evaluating experiment-guided ranking strategies. Though not perfectly accurate, its deviations resemble the noise observed in real wet-lab experiments and can foster more robust ranking strategies.
Figure 1: Overview of ranking strategies. Pre-experiment ranking is stateless and ignores feedback. Experiment-guided ranking with real experiments is stateful but infeasible to scale. Our simulator enables efficient development of ranking policies through simulated feedback before real deployment.
Building on this foundation, we develop an in-context reinforcement learning (ICRL) framework for experiment-guided hypothesis ranking. Within this framework, we instantiate a clustering-based agentic policy that decomposes hypotheses into functional components and groups them by shared mechanistic roles. After each experimental trial, the agent analyzes the tested hypothesis to infer which components contributed to its performance, then prioritizes untested hypotheses that incorporate the most promising functional elements. This enables efficient transfer of insights across structurally related candidates and helps navigate the hypothesis space more effectively. Experiments show that this policy significantly outperforms pre-experiment baselines and strong ablations. Combined with the simulator, our ICRL framework forms a general-purpose toolkit for studying experiment-guided ranking strategies, with our policy serving as a strong proof of concept. Figure 1 provides an overview of the three paradigms: pre-experiment ranking, experiment-guided ranking with real experiments, and our simulator-driven approach for developing ranking policies.
Overall, the contributions of this paper are:
@@
  • We formalize the task of experiment-guided ranking and highlight a key bottleneck in the natural sciences: the lack of scalable access to wet-lab experimental feedback. To address this, we propose the use of simulators and release a curated dataset of 124 scientific hypotheses with annotated performance collected from the literature.
@@
  • We introduce three conceptual foundations for simulating experimental feedback, formalize the simulation process mathematically, and construct a high-fidelity simulator that approximates real wet-lab outcomes under these foundations.
  • We present a clustering-based agentic ranking policy implemented within our ICRL framework. It generalizes from limited feedback and outperforms both pre-experiment baselines and ablation variants.
@@ 2 Methodology of Simulator Construction
2.1 Conceptual Foundations and Formalization
Our simulator construction is guided by three conceptual foundations—one assumption, one scientific principle, and one logical deduction—grounded in established principles of the natural sciences. Together, these provide a principled basis for modeling experimental outcomes of untested hypotheses, enabling systematic investigation of experiment-guided ranking strategies.
2.1.1 Conceptual Foundations
We posit that real experimental feedback within a hypothesis space can be simulated under the following conceptual foundations (A1–P1–D1):
@@ 2.1.1 Conceptual Foundations
    (A1: Local Optimum Assumption) A ground-truth hypothesis represents a dominant local optimum within its sufficiently local neighborhood of the hypothesis space.
@@
    (P1: Scientific Principle) Hypotheses that are more similar in their underlying structure or function tend to yield more similar experimental outcomes.
@@
    (D1: Logical Deduction) In practice, representations of hypothesis similarity are imperfect proxies, so the resulting performance landscape deviates from the ideal implied by A1 and P1, producing distortions such as noise, spurious local optima, or unexpected valleys.
A1, P1, and D1 are all reasonable and sufficiently grounded. P1 reflects the fundamental axiom that “structure determines properties, and properties determine outcome,” which underpins multiple disciplines: molecular structure and material function in Chemistry & Materials Science (Callister & Rethwisch, 1999; Hansch et al., 1995), crystal structure and physical properties in Physics (Wiley, 1986), and protein structure and biological function in Biology (Alberts et al., 2015). D1 follows logically from A1 and P1, since any practical representation of hypothesis similarity must introduce distortions. A1 is mostly valid but not guaranteed: even within a sufficiently small neighborhood, the labeled ground-truth hypothesis may not be the strict local optimum, as there could exist another hypothesis in that region with higher performance. This limitation, however, does not affect the simulator’s role in developing ranking policies, whose goal is to recover the labeled ground truth. When deployed in real experiments, any superior hypotheses beyond the labeled ground truth would be directly revealed, ensuring that policies developed under a simulator supported by A1 remain effective in practice.
(a) Idealized performance landscape (A1 + P1)).
(b) Realistic performance landscape (A1 + P1 + D1).
(c) Deviations from imperfect closeness estimation (D1).
Figure 2: Illustration of the three conceptual foundations (A1–P1–D1) for simulator construction.
Figure 2 visually illustrates these conceptual foundations. In the ideal scenario (Figure 2(a)), A1 ensures the presence of a dominant local optimum, while P1 enforces that hypotheses closer in structure or function to this optimum yield more similar outcomes. Together, these yield a smooth, unimodal performance landscape, where Euclidean distance in hypothesis space faithfully reflects structural and functional similarity. However, practical scenarios differ substantially, as the measured distance (“closeness”) between hypotheses—whether estimated by scientists or LLMs—may not faithfully capture structural and functional similarity. For instance, a chemical hypothesis might contain a useful functional component whose contribution is underrepresented, placing it farther from the dominant peak than warranted and creating a spurious secondary maximum. Conversely, a weaker hypothesis may appear deceptively close to the optimum, forming a local valley. These distortions yield a more irregular performance landscape, as illustrated in Figure 2(b), with unexpected secondary peaks and valleys. Figure 2(c) further isolates these deviations, highlighting the gap between the idealized oracle landscape and practical estimates of closeness.
We now formalize these foundations by defining a mathematical model that makes explicit the relationship between hypothesis embeddings, similarity, and performance.
@@ 2.1.2 Mathematical Formulation
$f(h,h^{*};q,\phi^{*}(\cdot))=\frac{1}{(2\pi\sigma^{2})^{d/2}}\exp(-\frac{\|\phi^{*}(h\mid q)-\phi^{*}(h^{*}\mid q)\|^{2}}{2\sigma^{2}}),$ (1)
@@
We model the idealized performance surface as a Gaussian-like function centered at $\phi^{*}(h^{*}\mid q)$, yielding a strictly unimodal landscape that decays smoothly with increasing distance from the optimum $h^{*}$ (Figure 2(a)). While the true performance landscape in feature space may not be strictly Gaussian, the isotropic Gaussian form serves as a tractable and interpretable approximation in the latent space. This modeling choice directly reflects A1 and P1.
@@
where $\phi(\cdot\mid q)$ is a practical embedding function that maps each hypothesis $h$ into (somewhat distorted) positions in the latent hypothesis space for a research question $q$, and $\epsilon(h\mid q)$ represents a systematic correction term that accounts for the discrepancy between oracle embedding $\phi^{*}(\cdot\mid q)$ and the practical embedding $\phi(\cdot\mid q)$ under the context of $q$. As a result, the practical embedding $\tilde{\mathcal{H}}$ introduces systematic distortions in the latent space, leading to spurious optima or valleys—effectively transforming the unimodal ideal surface into a noisier, multimodal one (Figure 2(b)).
@@
2.2 A Practical Implementation of $\phi(\cdot)$ with Prior Knowledge
@@
As discussed in § 2.1, the core objective of the simulator is to construct an embedding function $\phi(\cdot)$ that maps each hypothesis $h$ into a latent space such that distances in this space reflect meaningful functional differences. Through extensive discussions with domain experts, we observe that a scientific hypothesis succeeds in addressing a research question primarily due to its underlying mechanisms.
Specifically, an effective hypothesis typically comprises a set of scientifically meaningful components—each contributing to distinct yet complementary sub-mechanisms—which together enable the overall reaction to fulfill its intended function. The specific prompts and examples for extracting key components and inferring mechanisms are provided in § A.
@@
The simulation begins by decomposing both the candidate and ground truth hypotheses into a set of key functional components, and identifying the underlying mechanism associated with each component in the context of the research question. The decomposition of $h^{*}$ is performed first, serving as a reference. These reference components and mechanisms guide the decomposition of $h$, ensuring alignment in both granularity and mechanistic interpretation.
@@
$S(h\mid q;h^{*})=(\prod_{i\in\mathcal{C}}\mathbf{1}_{{s_{i}>0}})\cdot(\sum_{i=1}^{K}w_{i}\cdot s_{i}),\quad\text{where}\quad\sum_{i=1}^{K}w_{i}=1$ (3)
@@ 3 Methodology of Experiment-Guided Ranking
3.1 Task Formulation
Given a research question $q$, a set of candidate hypotheses $\mathcal{H}$ is formed by selecting hypotheses generated by existing scientific discovery systems (Yang et al., 2025) and ground-truth hypotheses from top-tier scientific journals reporting high-quality lab experiments. The goal of experiment-guided ranking is to identify the optimal hypothesis $h^{*}\in\mathcal{H}$ with the highest experimental performance using an experiment executor $E$. Formally, we define the experiment executor as a function: